OpenStudy (anonymous):

hi how to solve this indices question?? 2^p x 5 = 40

OpenStudy (anonymous):

is this \[2^p\times 5=40\]?

OpenStudy (anonymous):

YUP

OpenStudy (anonymous):

or \[2^{5p}=40\]?

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

if it is the first one divide by 5 to get \[2^p=4\] and then see that \[p=2\]

OpenStudy (anonymous):

how did you divide by 5?

OpenStudy (anonymous):

incorrectly i see

OpenStudy (anonymous):

\[\frac{40}{5}=8\] not 4

OpenStudy (anonymous):

so you really get \[2^p=8\] making \[p=3\]

OpenStudy (anonymous):

yesh but can you explain

OpenStudy (anonymous):

how you got rid of 5?

OpenStudy (anonymous):

is it just divide 40 by 5 by shifting 5 to the right side?

OpenStudy (anonymous):

sure. i will write it step by step \[2^p\times 5 = 40\] \[\frac{2^p\times 5}{5}=\frac{40}{5}\] \[2^p=8\] \[p=3\] since \[2^3=8\]

OpenStudy (anonymous):

GOSH thank you very much man.. I really really appreciate it :D

OpenStudy (anonymous):

divide both sides by 5 gives equivalent equation. yw